I need help setting up an equation to solve this problem. Using a calender, I quickly determined the answer was Saturday, however, I am having problems writing an equation to solve this problem.
What day would yesterday be if Friday was four days before the day after tommorrow?

Answer by Mathtut(3670)

let x be the unknown day of the week
:
if x+2-4=friday--->x-2=friday
:
we are looking for x-1
:
if you add 1 to both sides of the equation you get
:
x-2+1=friday +1
:
x-1=friday+1
:
x-1=Saturday
:
I dont know how complicated you want this formula but maybe this will help